OpenStudy (anonymous):

Two sides of an equilateral triangle have lengths of 2x+2 and -3+17. Which of 11-x or 5x-6 could be the length of the third side?

OpenStudy (anonymous):

i got to the part where i solved for x and it equals 3 but then you"re supposed to plug it into the one of the eqautions and didn't know what to do?

OpenStudy (anonymous):

If your problem is written correctly, then x is not equal to 3. That's your first challenge with this problem. Find the correct value of "x".

OpenStudy (anonymous):

well are you supposed to write your first step as 2x+2=3x+17

OpenStudy (anonymous):

That looks like a conflict with how you stated your problem. In your problem, you state that one side is -3+17, not 3x + 17. Which one is it supposed to be? It cannot be both.

OpenStudy (anonymous):

no, the problem says two sides of an equilateral triangle have lengths 2x+2 and -3x+17

OpenStudy (anonymous):

Now that you have the side stated correctly, yes, x equals 3. Now, just find 1) what the length of a side is by substituting "3" into 2x + 2 = L 2) determine which equation equals L when using x=3. It's either 11-x or 5x-6

OpenStudy (anonymous):

well when i did step 1 : 2x+2=L 2 times 3 +2 =L i got 6+2=L so then 8=L

OpenStudy (anonymous):

So far so good. The sides equal 8. That gets you halfway there to your answer. Now, do the second part I outlined above in my last post.

OpenStudy (anonymous):

okay so i followed step 2 in your last post and i used the equation 11-x and put in 11-3 and got 11-3=8 so is that the right answer since L=8

OpenStudy (anonymous):

Yes! Very good! Good job!

OpenStudy (anonymous):

thank you!(:

OpenStudy (anonymous):

uw!

OpenStudy (anonymous):

I have a new question, could you help?